Would someone please help explain what the question means by infinite solutions with the term of x. The answers I’ve gotten are 1.16( with a repeating line above 6) and 1.4.
Solve the system by any method. (If there is no solution, enter NO SOLUTION. If there are an infinite number of solutions, enter the general solution in terms of x, where x is any real number.)
6x − 8y = −4.2
3x + 2y = 6.3



Answer by math_helper(2461)


If the two equations happened to be linearly dependent, you'd really only have one equation with two unknowns. In that scenario there are an infinite number of solutions.

For example:
6x-8y = -4.2
12x-16y = -8.4
Here the 2nd equation is simply the first one multiplied by 2, hence they are linearly dependent, and in trying to solve, the best you can do is say y = (1/8) (4.2 + 6x), for all x.